# Super(ficial)-alignment: Strong Models May Deceive Weak Models in Weak-to-Strong Generalization

### v1

@@ 4.2 Results and Analysis
Figure 7: Results of deception scores on the reward modeling task.
We set the confidence threshold $T$ to identify the knowns and unknowns of a model to 0.75, but we also report the results under different thresholds in Appendix F to show that the patterns of weak-to-strong deception are independent to the choice of $T$. The conflict strength factor $\alpha$ in the explicit conflict setting is set to $0.5$.
@@
Despite the promising weak-to-strong generalization results, we can also observe a potential security issue of knowledge deception from the deception scores in Figure 7. (1) First, the weak-to-strong deception exists in both conflict settings. For example, all the deception scores are around or above 0.2, and some even exceed 0.5 especially when the weak model is GPT-2-Base. Our hypothesis is, in the reward modeling scenario, the strong model can obtain the predicted probability, which is also the confidence score, of the weak model on each sample. In such case, the strong model can explicitly perceive the knowledge boundary of the weak model, which creates the condition for the strong model to prioritize sacrificing the performance in the unknown area of the weak model for aligning with another target. (2) Second, the deception issue intensifies as the capability gap between the strong and weak models increases. As we can clearly see, under the same weak model, the deception score increases as the strong model becomes more powerful; under the same strong model, the deception score decreases as the weak teacher becomes stronger. There could be two reasons for this: (2.1) One is the Strong-Known area becomes larger as the strong model improves, leading to a greater probability of conflict tax occurring in this area. (2.2) The other reason is the stronger model spontaneously and increasingly tends to deceive the weak model in Weak-Unknown. We will delve into this discussion deeper in Section 5.3 through more visualizations. All in all, these findings reveal a great challenge for supervising LLMs as they are becoming more and more intelligent.
@@ 5.1 Weak-to-Strong Preference Alignment
The general procedure of weak-to-strong preference alignment is similar to that in the reward modeling scenario, but the major difference in this case is the strong model only receives and aligns with the final result of preference order that the weak model predicts for two completions within each sample. Please refer to Appendix A for the details.
@@ 5.2 Experimental Settings
Alignment Methods We mainly conduct experiments with the most recent offline preference optimization algorithm SimPO (Meng et al., 2024), due to its strengths of reference-free and being unbiased to the response length. We also perform experiments on DPO (Rafailov et al., 2024), the results are in Appendix E. We leave the exploration on the online preference optimization frameworks (Schulman et al., 2017) to future work.
Models Besides the GPT-2-series and OPT-series models, we further include a recent and advanced LLM Mistral-7B-v0.1 (Jiang et al., 2023) in this scenario for more comprehensive explorations. In SimPO, we can get the corresponding confidence of model $\boldsymbol{\theta}$ on $(x;y_{c},y_{r})$ as
$M_{\boldsymbol{\theta}}(x)=\text{Sigmoid}(\pi_{\boldsymbol{\theta}}(y_{c}|x)-\pi_{\boldsymbol{\theta}}(y_{r}|x)),$ (6)
where $\pi_{\boldsymbol{\theta}}(y|x)=\frac{1}{|y|}\sum\nolimits_{i=1}^{|y|}\log P_{\boldsymbol{\theta}}(y_{i}|x,y_{<i})$ is the normalized model logit of completion $y$. We can then follow Eq. (4) and Eq. (5) to calculate the test accuracy and deception score in this scenario.
Weak-to-Strong Objectives Here, we consider the same three weak-to-strong objectives as that in the reward modeling scenario. The detailed illustrations and mathematical forms are in Appendix B.2.
Training Details Please refer to Appendix C.
Figure 8: Deception scores of weak-to-strong experiments in the preference alignment scenario.
@@ 5.3 Results and Analysis
The confidence threshold $T$ and conflict strength factor $\alpha$ are 0.75 and 0.5, respectively. The results of deception scores under different $T$s are put in Appendix F, and we explore the effect of $\alpha$ on the severity of deception in Section 6.1. We put the detailed results and discussions regarding the weak-to-strong generalization performance in Appendix D, while here we mainly focus on the analysis of weak-to-strong deception issue.
@@
Regarding the results of deception scores shown in Figure 8, the main conclusions remain same as that in the reward modeling scenario. That is, the weak-to-strong deception issue exists in the preference alignment scenario, and the severity of the deception is positively correlated with the capability gap between weak and strong models. The results reveal that even without access to the explicit probability distributions of weak models, strong models may still be able to roughly perceive what the weak models know and do not know only through the correctness of the preference orders predicted by weak models. Compared to the analysis made in the reward modeling scenario, it means that the condition required for the deception phenomenon to occur could actually be more relaxed.
@@
We now attempt to make deeper explorations on the remaining question in Section 4.2 about the causes to the increasing deception scores as the model capability gap increases. As briefly discussed before, one reason could be the fact of larger Strong-Known area of a more advanced strong model, which correspondingly increases the probability of conflict tax occurring in $S_{k}\cap W_{uk}$; the other reason is stronger models increasingly tend to deceive weak models. To answer this question, we further visualize the dynamic changes of conflict tax across all four knowledge areas when the weak model is GPT-2-Large and the strong model varies among the remaining four larger models in the explicit conflict setting. In each case, we report both the deception score and the proportion of samples falling within $S_{k}\cap W_{uk}$ to the entire knowledge space (denoted as $|S_{k}\cap W_{uk}|/|S_{k}\cup S_{uk}|$). The results are in Figure LABEL:fig:_scatter_plots. The first conclusion we can draw is, as the strong model becomes more intelligent, there is a clear pattern that the misalignment (red crosses) gradually shifts towards Strong-Known and Weak-Unknown, leading to the more severe weak-to-strong deception phenomenon. Secondly, the growth rate of the deception score is much higher than that of $|S_{k}\cap W_{uk}|/|S_{k}\cup S_{uk}|$. This indicates that the growth of Strong-Known area only has limited contribution to the intensifying severity of deception phenomenon, and the primary cause is likely to be that stronger models tend to be more prone to deceive weak models in their unknown areas.
@@
Figure 13: Deception scores in weak-to-strong preference alignment with different conflict strength factor $\alpha$s.
6 Deeper Analysis
6.1 Stronger Conflicting Strength Generally Leads to More Severe Deception
In main experiments, we fix the conflict strength factor $\alpha$ in the explicit conflict setting to 0.5. Here, we conduct extra experiments with a smaller factor $\alpha=0.25$, and put the comparison results in Figure 13. As we can expect, the deception score consistently increases as the degree of conflict increases.
6.2 Discussions on Possible Countermeasures
@@
Considering the severe consequences that weak-to-strong deception may lead to, here, we make discussions on two possible ways to mitigate it. The following experiments are conducted in the preference alignment scenario
@@
6.2.1 Only Using High-Confidence Samples Cannot Mitigate Deception
@@
Based on the hypothesis we have made in Section 5.3, the reason why a strong model could deceive a weak model in preference alignment is because it is provided with both the correctly and wrongly predicted samples from the weak model. Thus, one possible solution to avoid deception may be only providing those high-confidence samples from the weak model for weak-to-strong alignment. We conduct experiments in the implicit conflict setting to validate this hypothesis, and the detailed experimental settings are in Appendix G. We provide the results when the weak model is GPT-2-Large in Figure LABEL:subfig:_high-confidence_deception_score as an illustrative example, and leave the full results in Appendix G. Unfortunately, we get the negative conclusion that only supervising with high-confidence samples cannot mitigate the deception phenomenon. This implies that there are deeper mechanisms to explain how strong models perceive the knowledge boundaries of weak models and exhibit deceptive behaviors. This can be an interesting direction for future work.
6.2.2 Bootstrapping Can Mitigate Deception to Some Extent
@@
Inspired by the relationship between deception severity and models’ capability gap, we are curious whether employing a bootstrapping method with an intermediate model would result in a lower deception score compared to directly using the weak model to supervise the strong model. In this case, we make the weak model first supervise an intermediate model and then let the intermediate model further supervise the ultimate strong model. We fix the ultimate strong model to Mistral-7B, and for each weak model, we select every model between it and Mistral as an intermediate model. Detailed settings are in Appendix H. The results are in Figure LABEL:subfig:_bootstrapping_accuracy and LABEL:subfig:_bootstrapping_deception_score. The results of cases when intermediate models are the same as weak models represent the results of directly using weak models to supervise Mistral. First, we can see that bootstrapping with an intermediate model can improve the generalization performance, which is consistent to the findings in Burns et al. (2023). More importantly, bootstrapping can indeed mitigate the deception issue to some extent, reflected in the consistent lower deception scores when intermediate models exist.
@@ 7 Conclusion and Discussion
We reveal and study a security issue in the weak-to-strong alignment, called the weak-to-strong deception. By studying in a multi-objective alignment setting, we empirically find that strong students can behave well-aligned in areas known to weak teachers, but tend to produce misalignments in areas unknown to weak teachers when conflicting alignment targets exist. Such a deception issue becomes more severe as the capability gap between weak and strong models increases. Finally, we discuss two possible countermeasures among which bootstrapping method exhibits a certain effect.
Regarding the future directions: (1) Given the fact that deception can occur even when only high-confidence samples from weak models are used for alignment, we need a more fundamental analysis of why and how strong models are prone to deceive weak models. (2) Given the limited effectiveness of bootstrapping, more effective mechanisms are needed to mitigate the deception issue.
Limitations
Though our study provides a comprehensive empirical analysis on the weak-to-strong deception issue, there are some limitations that can be interesting future work: (1) We mainly conduct experiments in the preference alignment scenario on two online preference optimization methods, SimPO (Meng et al., 2024) and DPO (Rafailov et al., 2024). Future work can explore the weak-to-strong deception issue on the offline preference optimization frameworks such as PPO (Schulman et al., 2017). (2) In our experiments, we only consider the case where the target alignment dimension is harmlessness, which is indeed an important alignment goal. However, there are some other dimensions that are also important for model alignment. For example, the deception issue also matters in the honesty alignment (Cheng et al., 2024), where the stronger model should not learn to deceive the weak model to intentionally make wrong responses on questions that the weak model does not know. (3) Due to the limited computational resources, we can only include models with 7B parameters as strong models. We have observed a pattern that the deception issue may become severe as strong models continue to evolve. We hope future work to validate this phenomenon on larger models.
Ethical Statement
@@
In this paper, we aim to reveal a potential security issue behind the current promising weak-to-strong generalization phenomenon. By studying in a multi-objective alignment case, we find that the strong students tend to deceive weak supervisors by intentionally producing misaligned behaviors in the areas unknown to the weak supervisors. Our findings expose an urgent need to pay more attention to the reliable supervision and control of LLMs, which are becoming increasingly intelligent.
We have also included some preliminary discussions on how to mitigate the deception problem in Section 6.2. However, the effectiveness of them is still limited, so we call for future studies to propose solutions that are more effective.
@@
Appendix A The Complete Procedure for Performing Weak-to-Strong Preference Alignment
@@
    Use ground truth preference data $D_{gt}$ and a preference optimization method to align weak and strong base models, obtain $\boldsymbol{\theta}_{w}^{gt}$ and $\boldsymbol{\theta}_{s}^{gt}$.
@@
    Use $\boldsymbol{\theta}_{w}^{gt}$ to make preference predictions on a held-out set and get $D_{weak}=\{(x;y_{c}^{w},y_{r}^{w})\}$, where $(y_{c}^{w},y_{r}^{w})$ is the preference order predicted by the weak model. Notice that this preference order may be different from the ground truth preference order $(y_{c}^{gt},y_{r}^{gt})$.
@@
    Use $D_{weak}$ (and other alignment targets if exist) to perform preference optimization on the strong base model to get the weak-to-strong model $\hat{\boldsymbol{\theta}}_{s}^{w}$ or $\boldsymbol{\theta}_{s}^{w}$.
Appendix B Concrete Mathematical Forms of All Weak-to-Strong Objectives
B.1 Reward Modeling Scenario
Here, we introduce the different weak-to-strong objectives in our experiments in detail. Besides the target alignment goal (harmlessness), we introduce two kinds of additional conflicting alignment goals for simulating the multi-objective alignment setting. (1) Explicit Conflict: The strong student will be given a direct reward/loss towards the opposite of the target dimension once it makes wrong predictions during training:
$\boldsymbol{\theta}_{s}^{w}=$ $\mathop{\arg\min}_{\boldsymbol{\theta}_{s}}\mathbb{E}_{x\sim D_{weak}}\big{[}\mathcal{L}_{CE}\big{(}M_{\boldsymbol{\theta}_{s}}(x),M_{\boldsymbol{\theta}_{w}^{gt}}(x)\big{)}$ (7)
$+\alpha\mathcal{L}_{CE}\big{(}M_{\boldsymbol{\theta}_{s}}(x),0\big{)}\cdot\mathbb{I}_{\{M_{\boldsymbol{\theta}_{s}}(x)<0.5\}}\big{]},$
where $\mathcal{L}_{CE}$ is the CrossEntropy Loss, $\mathbb{I}$ is the indicator function, $\alpha$ controls the conflict strength. (2) Implicit Conflict: The strong model needs to align with both the weak supervision on the harmless data and another supervision on the helpful data. Here, we introduce extra 4K ground truth helpful samples $D_{helpful}$ from HH-RLHF (Bai et al., 2022a) into the weak-to-strong process. In this case, the weak-to-strong objective can be written as:
$\boldsymbol{\theta}_{s}^{w}=$ $\mathop{\arg\min}_{\boldsymbol{\theta}_{s}}\big{[}\mathbb{E}_{x\sim D_{weak}}\mathcal{L}_{CE}\big{(}M_{\boldsymbol{\theta}_{s}}(x),M_{\boldsymbol{\theta}_{w}^{gt}}(x)\big{)}$ (8)
$+\mathbb{E}_{x\sim D_{helpful}}\mathcal{L}_{CE}\big{(}M_{\boldsymbol{\theta}_{s}}(x),1\big{)}\big{]}.$

### v2

@@ 4.2 Results and Analysis
We set the confidence threshold $T$ to identify the knowns and unknowns of a model to 0.75, but we also report the results under different thresholds in Appendix H to show that the patterns of weak-to-strong deception are independent to the choice of $T$. The conflict strength factor $\alpha$ in the explicit conflict setting is set to $0.5$.
@@
Despite the promising weak-to-strong generalization results, we can also observe a potential security issue of knowledge deception from the deception scores in Figure 9. (1) First, the weak-to-strong deception exists in both conflict settings. For example, all the deception scores are around or above 0.2, and some even exceed 0.5 especially when the weak model is GPT-2-Base. According to the definition of deception score, a positive deception score already indicates the existence of the deception phenomenon, as we never expect the conflict tax to occur in the area of Weak-Unknown and Strong-Known. Our hypothesis is, in the reward modeling scenario, the strong model can obtain the predicted probability, which is also the confidence score, of the weak model on each sample. In such case, the strong model can explicitly perceive the knowledge boundary of the weak model, which creates the condition for the strong model to prioritize sacrificing the performance in the unknown area of the weak model for aligning with another target. (2) Second, the deception issue intensifies as the capability gap between the strong and weak models increases. As we can clearly see, under the same weak model, the deception score increases as the strong model becomes more powerful; under the same strong model, the deception score decreases as the weak teacher becomes stronger. There could be two reasons for this: (2.1) One is the Strong-Known area becomes larger as the strong model improves, leading to a greater probability of conflict tax occurring in this area. (2.2) The other reason is the stronger model spontaneously and increasingly tends to deceive the weak model in Weak-Unknown. We will delve into this discussion deeper in Section 5.3 through more visualizations. All in all, these findings reveal a great challenge for supervising LLMs as they are becoming more and more intelligent.
Figure 8: Deception scores on the reward modeling task.
Figure 9: Deception scores in the preference alignment scenario.
@@ 5.1 Weak-to-Strong Preference Alignment
The general procedure of weak-to-strong preference alignment is similar to that in the reward modeling scenario, but the major difference in this case is the strong model only receives and aligns with the final result of preference order that the weak model predicts for two completions within each sample. Please refer to Appendix B for the details.
@@ 5.2 Experimental Settings
Alignment Methods We mainly conduct experiments with the most recent offline preference optimization algorithm SimPO (Meng et al., 2024), due to its strengths of reference-free and being unbiased to the response length. We also perform experiments on DPO (Rafailov et al., 2024). We put the detailed experimental settings and full results on DPO in Appendix 31. We leave the exploration on the online preference optimization frameworks (Schulman et al., 2017) to future work.
Models Besides the GPT-2-series and OPT-series models, we further include a recent and advanced LLM Mistral-7B-v0.1 (Jiang et al., 2023) in main experiments in this scenario for more comprehensive explorations. We also conduct supplemental experiments on LLaMA-3-8B/70B (MetaAI, 2024a) and LLaMA-3.1-8B (MetaAI, 2024b) models to explore the weak-to-strong issue on larger models or same size models with more powerful capabilities. We put the detailed experimental settings, full results and discussion in Appendix G, while the main conclusions remain the same. In SimPO, we can get the corresponding confidence of model $\bm{\theta}$ on $(x;y_{c},y_{r})$ as
$M_{\bm{\theta}}(x)=\text{Sigmoid}(\pi_{\bm{\theta}}(y_{c}|x)-\pi_{\bm{\theta}}(y_{r}|x)),$ (9)
where $\pi_{\bm{\theta}}(y|x)=\frac{1}{|y|}\sum\nolimits_{i=1}^{|y|}\log P_{\bm{\theta}}(y_{i}|x,y_{<i})$ is the normalized model logit of completion $y$. We can then follow Eq. (7) and Eq. (8) to calculate the test accuracy and deception score in this scenario.
Weak-to-Strong Objectives Here, we consider the same three weak-to-strong objectives as that in the reward modeling scenario. Detailed illustrations and mathematical forms are in Appendix C.2.
Training Details Please refer to Appendix D.
@@ 5.3 Results and Analysis
The confidence threshold $T$ and conflict strength factor $\alpha$ are 0.75 and 0.5, respectively. The results of deception scores under different $T$s are put in Appendix H, and we explore the effect of $\alpha$ on the severity of deception in Figure 15. We put the detailed results and discussions regarding the weak-to-strong generalization performance in Appendix E, while here we mainly focus on the analysis of weak-to-strong deception issue.
@@
Regarding the results of deception scores shown in Figure 9, the main conclusions remain same as that in the reward modeling scenario. That is, the weak-to-strong deception issue exists in the preference alignment scenario, and the severity of the deception is positively correlated with the capability gap between weak and strong models. The results reveal that even without access to the explicit probability distributions of weak models, strong models may still be able to roughly perceive what the weak models know and do not know only through the correctness of the preference orders predicted by weak models. Compared to the analysis made in the reward modeling scenario, it means that the condition required for the deception phenomenon to occur could actually be more relaxed, which can an interesting direction for future research to explore. Furthermore, results on LLaMA-3/LLaMA-3.1-series models in Figure LABEL:fig:_experiments_on_llama3 re-validates that the essential factor that affects the deception severity is not solely the model scale, but the model capability. Finally, the comparison results of setting different conflict strength factors $\alpha$ in Figure 15 show that the deception score consistently increases as the degree of conflict increases.
@@
We now attempt to make deeper explorations on the remaining question in Section 4.2 about the causes to the increasing deception scores as the model capability gap increases. As briefly discussed before, one reason could be the fact of larger Strong-Known area of a more advanced strong model, which correspondingly increases the probability of conflict tax occurring in $S_{k}\cap W_{uk}$; the other reason is stronger models increasingly tend to deceive weak models. To answer this question, we further visualize the dynamic changes of conflict tax across all four knowledge areas when the weak model is GPT-2-Large and the strong model varies among the remaining four larger models in the explicit conflict setting. The results are in Figure LABEL:fig:_scatter_plots. Also, in each case, we report both the deception score and the proportion of samples falling within $S_{k}\cap W_{uk}$ to the entire knowledge space (denoted as $|S_{k}\cap W_{uk}|/|S_{k}\cup S_{uk}|$). The results are in Figure 15. The first conclusion we can draw from Figure LABEL:fig:_scatter_plots is, as the strong model becomes more intelligent, there is a clear pattern that the misalignment (red crosses) gradually shifts towards Strong-Known and Weak-Unknown, leading to the more severe weak-to-strong deception phenomenon. Secondly, according to Figure 15, the growth rate of the deception score is much higher than that of $|S_{k}\cap W_{uk}|/|S_{k}\cup S_{uk}|$ (this pattern also consistently exists in other experiments with different weak models and types of conflict). This indicates that the growth of Strong-Known area only has limited contribution to the intensifying severity of deception phenomenon, and the primary cause is likely to be that stronger models themselves tend to be more prone to deceiving weak models in weak model’s unknown areas.
We also conduct a case study in Table 1 to provide a concrete example belonging to the weak-to-strong deception phenomenon, please refer to Appendix K for the detailed discussion.
Figure 14: The comparison between the increasing trends of deception score and proportion of the Strong-Known and Weak-Unknown area.
@@
Figure 15: Deception scores in weak-to-strong preference alignment with different explicit conflict strength factor $\alpha$s.
6 Discussions on Possible Countermeasures
@@
Considering the severe consequences that weak-to-strong deception may lead to, here, we make discussions on two possible ways to mitigate it. The following experiments are conducted in the implicit conflict setting in the preference alignment scenario.
@@
6.1 Only Using Correct High-Confidence Samples Cannot Mitigate Deception
@@
Based on the hypothesis we have made in Section 5.3, the reason why a strong model could deceive a weak model in preference alignment is might be that it is provided with both the correctly and wrongly predicted samples from the weak model. Thus, one possible solution to avoid deception may be only providing those correct high-confidence samples from the weak model for weak-to-strong alignment. We conduct experiments in the implicit conflict setting to validate this hypothesis, and the detailed experimental settings are in Appendix I. We provide the results when the weak model is GPT-2-Large in Figure LABEL:subfig:_high-confidence_deception_score as an illustrative example, and leave the full results in Appendix I. Unfortunately, we get the negative conclusion that only supervising with high-confidence samples cannot mitigate the deception phenomenon. This implies that there are deeper mechanisms to explain how strong models perceive the knowledge boundaries of weak models and exhibit deceptive behaviors, which can be an interesting direction for future work. For example, strong models may possibly infer the areas of knowledge where teachers excel and struggle based on the portion and distribution of samples that teachers provide across different domains in this setting.
6.2 Bootstrapping Can Mitigate Deception to Some Extent
@@
Inspired by the relationship between deception severity and models’ capability gap, we are curious whether employing a bootstrapping method with an intermediate model (Burns et al., 2024) would result in a lower deception score compared to directly using the weak model to supervise the strong model. In this case, we make the weak model first supervise an intermediate model and then let the intermediate model further supervise the ultimate strong model. We fix the ultimate strong model to Mistral-7B, and for each weak model, we select every model between it and Mistral as an intermediate model. Detailed experimental settings are in Appendix J. The results are in Figure LABEL:subfig:_bootstrapping_accuracy and LABEL:subfig:_bootstrapping_deception_score. The results of cases when intermediate models are the same as weak models represent the results of directly using weak models to supervise Mistral-7B. Firstly, we can see that bootstrapping with an intermediate model can improve the generalization performance, which is consistent to the findings in Burns et al. (2024). More importantly, bootstrapping can indeed mitigate the deception issue to some extent, reflected in the consistently lower deception scores when intermediate models exist. The reason could be that some cases originally unknown to the weak model are now known to the intermediate model, making it difficult for the strong model to deceive in those cases.
@@ 7 Conclusion and Discussion
In this paper, we reveal and study a security issue in the weak-to-strong alignment, called the weak-to-strong deception. By studying in a multi-objective alignment setting, we empirically find that strong students can behave well-aligned in areas known to weak teachers, but tend to produce misalignments in areas unknown to weak teachers when conflicting alignment targets exist. Such a deception issue becomes more severe as the capability gap between weak and strong models increases, which introduces a greater challenge for humans to reliably supervise super AI as it continuously becomes smarter and more intelligent in the future. Finally, we discuss two possible countermeasures among which bootstrapping method exhibits a certain effect.
Regarding the future directions: (1) We have mainly conducted a preliminary study on the issue of weak-to-strong deception in a specific multi-objective alignment scenario. Future work could make more explorations in other scenarios, such as the spontaneous deception by AI (Denison et al., 2024). (2) Given our concerning experimental findings, we need a more fundamental analysis of why and how strong models are prone to deceiving weak models. (3) Given the limited effectiveness of bootstrapping method, more effective mechanisms are needed to mitigate the deception issue.
Ethics Statement
@@
In this paper, we aim to reveal a potential security issue behind the current promising weak-to-strong generalization phenomenon. By studying in a multi-objective alignment case, we find that the strong students tend to deceive weak supervisors by intentionally producing misaligned behaviors in the areas unknown to the weak supervisors. Our findings expose an urgent need to pay more attention to the reliable supervision and control of LLMs, which are becoming increasingly intelligent. We have also included some preliminary discussions on how to mitigate the deception problem in Section 6. However, the effectiveness of them is still limited, so we call for future studies to propose solutions that are more effective.
Appendix A Limitations
Though our study provides a comprehensive empirical analysis on the weak-to-strong deception issue, there are some limitations that can be interesting future work: (1) We mainly conduct experiments in the preference alignment scenario on two offline preference optimization methods, SimPO (Meng et al., 2024) and DPO (Rafailov et al., 2024). Future work can explore the weak-to-strong deception issue on the online preference optimization frameworks such as PPO (Schulman et al., 2017). (2) In our experiments, we only consider the case where the target alignment dimension is harmlessness, which is indeed an important alignment goal. However, there are some other dimensions that are also important for model alignment. For example, the deception issue also matters in the honesty alignment (Cheng et al., 2024), where the stronger model should not learn to deceive the weak model to intentionally make wrong responses on questions that the weak model does not know.
@@ Appendix A Limitations
Appendix B The Complete Procedure for Performing Weak-to-Strong Preference Alignment
@@
    Use ground truth preference data $D_{gt}$ and a preference optimization method to align weak and strong base models, obtain $\bm{\theta}_{w}^{gt}$ and $\bm{\theta}_{s}^{gt}$.
@@
    Use $\bm{\theta}_{w}^{gt}$ to make preference predictions on a held-out set and get $D_{weak}=\{(x;y_{c}^{w},y_{r}^{w})\}$, where $(y_{c}^{w},y_{r}^{w})$ is the preference order predicted by the weak model. Notice that this preference order may be different from the ground truth preference order $(y_{c}^{gt},y_{r}^{gt})$.
@@
    Use $D_{weak}$ (and other alignment targets if exist) to perform preference optimization on the strong base model to get the weak-to-strong model $\tilde{\bm{\theta}}_{s}^{w}$ or $\bm{\theta}_{s}^{w}$.
Appendix C Concrete Mathematical Forms of All Weak-to-Strong Objectives
C.1 Reward Modeling Scenario
Here, we introduce the different weak-to-strong objectives in our experiments in detail. Besides the target alignment goal (harmlessness), we introduce two kinds of additional conflicting alignment goals for simulating the multi-objective alignment setting.
(1) Explicit Conflict: The strong student will be given a direct reward/loss towards the opposite of the target dimension once it makes wrong predictions during training:
$\bm{\theta}_{s}^{w}=$ $\mathop{\arg\min}_{\bm{\theta}_{s}}\mathbb{E}_{x\sim D_{weak}}\big{[}\mathcal{L}_{CE}\big{(}M_{\bm{\theta}_{s}}(x),M_{\bm{\theta}_{w}^{gt}}(x)\big{)}+\alpha\mathcal{L}_{CE}\big{(}M_{\bm{\theta}_{s}}(x),0\big{)}\cdot\mathbb{I}_{\{M_{\bm{\theta}_{s}}(x)<0.5\}}\big{]},$ (10)
where $\mathcal{L}_{CE}$ is the CrossEntropy Loss, $\mathbb{I}$ is the indicator function, $\alpha$ controls the conflict strength.
(2) Implicit Conflict: The strong model needs to align with both the weak supervision on the harmless data and another supervision on the helpful data. Here, we introduce extra 4K ground truth helpful samples $D_{helpful}$ from HH-RLHF (Bai et al., 2022a) into the weak-to-strong process. In this case, the weak-to-strong objective can be written as:
$\bm{\theta}_{s}^{w}=$ $\mathop{\arg\min}_{\bm{\theta}_{s}}\big{[}\mathbb{E}_{x\sim D_{weak}}\mathcal{L}_{CE}\big{(}M_{\bm{\theta}_{s}}(x),M_{\bm{\theta}_{w}^{gt}}(x)\big{)}+\mathbb{E}_{x\sim D_{helpful}}\mathcal{L}_{CE}\big{(}M_{\bm{\theta}_{s}}(x),1\big{)}\big{]}.$ (11)